OpenStudy (anonymous):

The ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6.

OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? Answer (1, 7) (5, 0) (2, -1) (0, 5)

OpenStudy (compassionate):

Hello, in order to find out if you have a solution the a system of equations, you must take each term and plug it in to the slope-intercept equations and see if the end result is true or false. I will do the first one for you Lets plug in (1, 7) x = 1, y = 7 y = -3x + 5 7 = -3(1) + 5 7 = -3 + 5 7 = 2 This is false because seven does not equal two. x - y = 6 -y = -x + 6 y = x - 6 7 = 1 - 6 7 = -5 This is also false. 7 = 7, not -5. Do the same thing for the rest of the points.

OpenStudy (anonymous):

Thanks! could you help me with this one to please? Is the system consistent or inconsistent? y = -1/2x - 1 and y = 3x + 3

OpenStudy (compassionate):

Sorry - I am not familiar with those type of mathematics.

OpenStudy (anonymous):

Which system has a solution in which the x value is greater than one? y = -2/3 x and y = 3x - 3 y = 3x + 2 and y = -1/4 x + 1 y = -1/4 x + 1 and y = -2/3 x y = 3x + 2 and y = -2/3 x